If a farsighted person has a near point that is 0.600 m from the eye, what is the focal length f2 of the contact lenses that the person would need to be able to read a book held at 0.350 m from the person's eyes?
Answer #1

The focal length of the contact lens is

\frac{1}{f} = \frac{1}{d_i}+\frac{1}{d_o}

\frac{1}{f} = \frac{1}{0.35}+\frac{1}{-0.6}

f = 0.84m